OpenStudy (anonymous):

Correct me if I am wrong. Solve the quadratic equation m2+ 2m - 24 = 0 by factoring. Show all your steps.

OpenStudy (anonymous):

We put this like m^2 + bm - c then we find what two numbers multiplied together = b and added together = c

OpenStudy (anonymous):

So, -4 x 6 = -24 and -4 + 6 = 2

OpenStudy (anonymous):

And I am not sure if there is anymore to the equation or not.

OpenStudy (anonymous):

you are on the right track, just not quite done.

OpenStudy (anonymous):

you found two factors that multiply to -24 and add to 2.... -4 and 6 So the factored form of m^2 + 2m - 24 = 0 looks like: (m - 4) (m + 6) = 0 Because... if you multiply those parenthesis out, you get m^2 -4m +6m - 24 = 0 and the -4m and +6m simplify back to +2m like in the original.

OpenStudy (anonymous):

Now that you have the factored form as (m - 4) (m + 6) = 0 you need to solve for the m values that make this true. It will be true when the left side of the equation is = 0. That happens when EITHER parenthesis term is equal 0. So m = 4 will make the first term 0. And m = -6 will make the 2nd term 0. That means that m = 4 and m = -6 are solutions to your original equation. You have "solved the quadratic equation by factoring" :)

OpenStudy (anonymous):

Ahhh, I see what you did there. Thank you sir! (:

OpenStudy (anonymous):

The whole reason to find those factors that multiply to be the last number and add to be the coefficient of middle term is to set up that final factored form of the equation... then you can get the solution values for m. Glad to help! Good luck!